OpenStudy (faman39):

A spinner is divided into 8 equal segments. The segments are labeled with the letters A, B, E, I, L, O, T, U. Match each compound event with its probability. Round to the nearest percent, if necessary. Matching : a. 16% b. 5% c. 14% d. 39% P ( vowel and vowel) P ( consonant and T) P ( consonant and consonant) P ( letter before D and a letter after H)

OpenStudy (faman39):

@blue can u able to help me out with this?

OpenStudy (faman39):

@cwrw238 can u able to help me out this?

OpenStudy (faman39):

@ksaimouli can u able to help me with this?

OpenStudy (faman39):

@Zarkon can u able to help me?

OpenStudy (faman39):

@Agent_Sniffles can u able to help me with it?

OpenStudy (kropot72):

The total sample space for two spins is 8 * 8 possible outcomes. The sample space for vowel and vowel is 5 * 5 possible outcomes. \[P(Vowel+Vowel)=\frac{5\times 5}{8\times 8}=?\]

OpenStudy (kropot72):

The probability of spinning a consonant is 3/8. The probability of spinning a T is 1/8. P(consonant and T) = 3/8 * 1/8 = ?